A Bibliography of Volumes in Lecture Notes in Computer Science (2012)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

             E-mail: beebe@math.utah.edu, beebe@acm.org,
                       beebe@computer.org (Internet)
           WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.03

Title word cross-reference

3 [DHtM^{+}12, YYH^{+}12].

10th [CVTP12, FB12, FN12, GVB12, Jón12a, Jón12b, KMMT12, TCYZ12]. 11-14 [HGPH12]. 11th [BKL^{+}12b, BKL^{+}12a, MPT12, RKS^{+}12a, RKS^{+}12b, ST12a]. 12-14th [TR12]. 12th [GPR^{+}12, HM12, JY12]. 13th [BDDM12, DLW12, Gel12a, Gel12b, KR12, MDPQA12a, MDPQA12b, TR12]. 14th [RZ12, SWJX12, TR12]. 15th [AJN12, Bir12, Dan12, FBM12, MSK^{+}12, OO12, dLZ12]. 16th [Cho12, Lau12, Sch12, TCHB12b, TCHB12a]. 17th [gLPZ^{+}12b, gLPZ^{+}12a, YYH^{+}12].

18th [BV12, FK12, MV12, RD12, SMH^{+}12]. 19th [vKS12].

2008-2010 [MN12]. 20th [MK12]. 21st [O’B12, Sei12]. 25th [HRB12]. 27th [Mac12]. 28-30th [CSA^{+}12b, CSA^{+}12a].

30th [CSA^{+}12b, CSA^{+}12a]. 31- [HNP12]. 31st [PJ12]. 34th [BYdVZ^{+}12]. 38th [BFG^{+}12].

4th [AS12, BLS12, GANACBdCdV12, GAL12, GP12, HNP12, IS12, JMP12, PCN12c, PCN12a, PCN12b, PSD12, SA12, SSZ12, VMB^{+}12].

5th [Ho12b, Ho12a, YGE12].

60th [ACÇ^{+}12, CS12, DKN12, DKS12]. 65th [Nac12]. 6th [DMV12, GANACBdCdV12, KH12, MR12, RiN12].

7th [BL12, BCC^{+}12, CBS^{+}12, CSA^{+}12b, CSA^{+}12a, ENO12, HGBF12, HGGG12, HGPH12, JP12, KBC^{+}12, LCB12, LS12, MFG12, PRK12].

8th [BMSW12, CP12, CCAS12, CVV12, CDN12, PNPP12, RR12, RSW12, VAR12].

9th [ACL12, AdBB12, BJP12, Cra12, LST^{+}12, PH12, SSVW12, SH12a, SOP12].

A4MMC [VMvN12]. AAIM [SLSW12]. AAMAS [DHtM^{+}12, VKG12]. Abdominal [YSL12]. Abstract [KR12]. Achievements [BSW12]. ACIIDS [PCN12c, PCN12a, PCN12b]. Action [ACÇ^{+}12, HS12]. Actions [vEV12]. Active [TR12]. Ad [ENO12]. Adaptive [VKG12]. ADCONS [TPCB12]. ADMI [CBS^{+}12]. Advanced [gLPZ^{+}12b, gLPZ^{+}12a, YYH^{+}12, DHtM^{+}12, HGBF12, HGGG12, TPCB12]. Advances [GTP12, PRK12, SH12a, AK12, BYdVZ^{+}12, Ho12b, Ho12a, LCB12, PJ12, SMH^{+}12, SSZ12, TCHB12b, TCHB12a]. Age [WJL^{+}12]. Agent [CDN12, DHtM^{+}12, DLW12, SSVW12, VSMS12]. Agents [CBS^{+}12, VKG12, CBS^{+}12]. AI [BJP12]. Aided [MDPQA12a, MDPQA12b]. AIME [RtTM12]. ALA [VKG12]. Algebraic [MK12, HNP12]. Algorithmic [SLSW12]. Algorithmics [SLSW12]. Algorithms [ENO12, FK12, RiN12, SOP12, ENO12]. ALGOSENSORS [ENO12]. AMAS [VMvN12]. AMAS-BT [VMvN12]. American [FB12]. Amherst [JP12]. AMPLE [DHtM^{+}12]. Analysis [FK12, IS12, LSW12, PSD12, SH12b, TCYZ12]. ANB [HNP12]. Annual [ACL12, Cho12, PJ12]. AOSE [DHtM^{+}12].

Applications [ACL12, AJN12, BMSW12, CCAS12, DMV12, HGGG12, HGPH12, JY12, Lau12, gLPZ^{+}12b, gLPZ^{+}12a, MP12, MOT12, Nac12, PNPP12, PJ12, SWJX12, WZH^{+}12, YSL12, YYH^{+}12, CAC^{+}12]. Applied [CHB^{+}12, Jón12a, Jón12b]. Approaches [dLZ12]. Approximation [SOP12]. April [AS12, BYdVZ^{+}12, Bir12, CVTP12, CAC^{+}12, Cho12, DG12, FB12, FK12, GVB12, GP12, HM12, HLKX12, gLPZ^{+}12b, gLPZ^{+}12a, LST^{+}12, MRC12, MPT12, MP12, MSK^{+}12, O’B12, PJ12, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, RSW12, Sei12, SWJX12, VMB^{+}12, WZH^{+}12, YGE12, YYH^{+}12, dLZ12]. APWeb [SWJX12, WZH^{+}12]. ARC [CCAS12]. Architecture [VMvN12, HRB12]. Architectures [CCAS12]. ARCS [HRB12]. Areas [MV12]. Arequipa [FB12]. Argumentation [MOT12]. ARMS [DHtM^{+}12]. Art [MRC12]. Artificial [BV12, CSA^{+}12b, CSA^{+}12a, HGGG12, RKS^{+}12a, RKS^{+}12b]. Asia [SWJX12, TCHB12b, TCHB12a]. Asia-Pacific [SWJX12]. Asian [PCN12c, PCN12a, PCN12b]. Aspects [BL12, HGGG12, KKW12, RZ12, SLSW12, vEV12]. Assembly [ÁCD^{+}12]. Athens [SH12a]. Atlases [CKP^{+}12]. Auckland [DKN12]. August [ADB^{+}12b, ADB^{+}12a, GPR^{+}12, HNP12, HGBF12, HGGG12, HGPH12, JY12, MV12]. Austria [AdBB12, HNP12, PSD12, SMH^{+}12, TR12]. Automata [DMV12]. Autonomic [GTP12]. Autonomous [ENO12, GANACBdCdV12]. Avignon [BCC^{+}12]. Aviv [TCYZ12].

Bangladesh [RiN12]. Barcelona [BYdVZ^{+}12, Cho12, MOT12]. Based [MGSM12, IS12, VSMS12]. Bay [DDS12]. Behavioral [YGE12]. Beijing [ACL12, HLKX12, IS12, SLSW12]. Belgium [GANACBdCdV12, PNPP12, VW12].

Bernhard [DKS12]. Beyond [DKN12]. Bhubaneswar [RR12]. Bio [HGPH12]. Bio-Inspired [HGPH12]. Bioinformatics [GVB12, BMSW12, FN12]. Biologically [MRC12]. Biology [Cho12, HNP12]. Birthday [DKN12, ACÇ^{+}12, CS12, DKS12, Nac12]. Bled [RtTM12]. Block [CWP12]. BNCOD [Mac12]. Bordeaux [ADB^{+}12b, ADB^{+}12a]. Braga [SA12]. British [Mac12]. BT [VMvN12]. Building [BSW12]. Busan [gLPZ^{+}12b, gLPZ^{+}12a, YYH^{+}12].

CA [Dun12, KS12]. Calude [DKN12]. Cambridge [LST^{+}12, PJ12]. Camera [IS12]. Camera-Based [IS12]. Canada [CKP^{+}12, CDN12, GAL12, MV12, MGSM12, YSL12]. Canada-France [GAL12]. Canaria [MDPQA12a, MDPQA12b]. Care [RtTM12]. CBDAR [IS12]. CC [O’B12]. CCPI [ADB^{+}12b, ADB^{+}12a]. CDS [MGSM12]. CEE [SSZ12]. CEE-SET [SSZ12]. Cells [LST^{+}12]. Centered [HKW12]. Central [SSZ12]. CGWS [ADB^{+}12b, ADB^{+}12a]. Challenge [KKW12]. Challenges [CKP^{+}12]. Channel [SH12b]. Checking [KR12]. China [ACL12, BDDM12, CHB^{+}12, CCAS12, HLKX12, HGBF12, HGGG12, HGPH12, IS12, LCB12, RSW12, SWJX12, SLSW12, WJL^{+}12, WZH^{+}12]. CICLing [Gel12a, Gel12b]. Ciphers [CWP12]. Clermont [CP12]. Clermont-Ferrand [CP12]. Clinical [MGSM12, YSL12]. Clouds [VAR12]. CMC [GPR^{+}12]. Coimbra [CVTP12]. College [YGE12]. Colloquium [RKM^{+}12]. Combinatorial [BJP12, HM12]. Communication [KMMT12, VMB^{+}12]. Compiler [O’B12]. Complexity [BW12]. CompNets [BBK12]. Component [BL12]. Components [AdBB12]. Computation [ACL12, CAC^{+}12, GVB12, HM12, MR12, RiN12, RKS^{+}12c, VW12, DKN12].

Computational [ABH12, Bir12, CKP^{+}12, Cho12, GTP12, YSL12, CVTP12, Gel12a, Gel12b]. Computer [BFG^{+}12, BtHRBB12, DKN12, KBC^{+}12, MDPQA12a, MDPQA12b, VMvN12]. Computing [BDDM12, CCAS12, GTP12, GPR^{+}12, HRB12, HGBF12, HGGG12, HGPH12, Jón12a, Jón12b, LCB12, RR12, RKM^{+}12, RKS^{+}12a, RKS^{+}12b, Sch12, TPCB12, YGE12, KKW12]. Conceptual [DKS12]. Concurrency [JDK12]. Conference [ABH12, ACL12, AS12, AJN12, BYdVZ^{+}12, BJP12, BKL^{+}12b, BKL^{+}12a, BFG^{+}12, Bir12, BV12, BCC^{+}12, BDDM12, BtHRBB12, CVTP12, Cho12, CVV12, CSA^{+}12b, CSA^{+}12a, Cra12, Dan12, DMV12, DG12, DLW12, Dun12, FBM12, FK12, Gel12a, Gel12b, GPR^{+}12, GVB12, HM12, HLKX12, HRB12, HNP12, HGBF12, HGGG12, HGPH12, Jón12a, Jón12b, JMP12, KA12, KS12, KMMT12, KR12, KMMP12, Lau12, gLPZ^{+}12b, gLPZ^{+}12a, LCB12, LST^{+}12, Mac12, MRC12, MSK^{+}12, MDPQA12a, MDPQA12b, O’B12, PCN12c, PCN12a, PCN12b, PRK12, PH12, PJ12, RR12, RD12, RKS^{+}12a, RKS^{+}12b, RSW12, Sch12, SMH^{+}12, SWJX12, SA12, SLSW12, SSZ12, TR12, TCHB12b, TCHB12a, TCYZ12, TPCB12, YGE12, YYH^{+}12, dLZ12]. Conferences [Bir12, BKL^{+}12c, DG12, FK12, O’B12, Sei12, dLZ12]. Conjunction [CKP^{+}12, RKM^{+}12, RKS^{+}12c, YSL12]. Construction [FK12, O’B12]. Constructive [SH12b]. Content [MGSM12]. Content-Based [MGSM12]. Contraint [BJP12]. Copenhagen [MFG12]. Correlation [BW12]. Coruña [DMV12]. COSADE [SH12b]. COST [HS12]. CPAIOR [BJP12]. Cristian [DKN12]. Critical [LSW12]. Cryptanalysis [BW12, CWP12]. Cryptographer [Dun12]. Cryptographic [PJ12]. Cryptography [Cra12, DDS12, Dan12, FBM12, MV12, Nac12]. Cryptology [Dun12, PJ12]. CT [Dun12]. CT-RSA [Dun12]. Cultural [YGE12]. Current [BFG^{+}12, HK12]. Curves [BCC^{+}12]. Cyprus [HK12, VAR12]. Czech [BBK12, BKL^{+}12b, BKL^{+}12a, BFG^{+}12, KBC^{+}12].

Dallas [BMSW12]. DALT [SSVW12]. Darmstadt [FBM12, SH12b]. DASFAA [gLPZ^{+}12b, gLPZ^{+}12a, YYH^{+}12]. Data [BW12, CBS^{+}12, CHB^{+}12, DDS12, Dan12, GVB12, HKW12, Mac12, Shi12, TCHB12b, TCHB12a, GANACBdCdV12, KA12, Mac12]. Data- [HKW12]. Database [PCN12c, PCN12a, PCN12b, gLPZ^{+}12b, gLPZ^{+}12a, YYH^{+}12]. Databases [Mac12]. December [AdBB12, TPCB12, VAR12]. Decision [MGSM12]. Declarative [RZ12, SSVW12]. Dedicated [CS12, DKN12, DKS12, Nac12]. Defense [LSW12]. Delft [KH12]. Delhi [Gel12a, Gel12b]. Denmark [MFG12]. Dependability [Sch12]. Design [MRC12, SH12b, PRK12]. DESRIST [PRK12]. Development [MK12]. Dexter [CS12]. DFT [Sch12]. Dhaka [RiN12]. Differential [CWP12]. Discovery [LV12, TCHB12b, TCHB12a]. Distributed [BSW12, BDDM12, RR12]. DOCM [DHtM^{+}12]. Doctoral [HK12, KBC^{+}12, RKM^{+}12]. Document [IS12]. DPM [GANACBdCdV12]. DQDI [YYH^{+}12]. Drawing [vKS12]. Dundee [Mac12].

e-Infrastructure-PL-Grid [BSW12]. EAMA [VMvN12]. East [SSZ12]. EC [RKS^{+}12c]. ECIR [BYdVZ^{+}12]. Econ [HS12]. Economics [HS12, VAR12]. Edutainment [PCM^{+}12]. Ein-Gedi [BtHRBB12]. Eindhoven [BLS12, vKS12]. Elba [MN12]. Empirical [MN12]. Engineering [AS12, GTP12, HK12, KA12, Kie12, KBC^{+}12, MN12, RD12, SA12, SSZ12, dLZ12, BLS12]. Entities [ENO12]. Ershov [CVV12]. ESOP [Sei12]. Espoo [AJN12]. Essays [ACÇ^{+}12, CS12, DKS12, Nac12]. Essen [RD12]. ESSoS [BLS12]. Estonia [Bir12, DG12, FK12, Lau12, O’B12, Sei12, dLZ12]. ESWC [GCFA12]. ETAPS [Bir12, DG12, FK12, O’B12, Sei12, dLZ12]. Etelsen [MK12]. ETICS [BBK12]. Euro [ADB^{+}12b, ADB^{+}12a]. Euro-Par [ADB^{+}12b, ADB^{+}12a]. EUROCAST [MDPQA12a, MDPQA12b]. EUROCRYPT [PJ12]. EuroGP [MSK^{+}12]. European [BYdVZ^{+}12, Bir12, DG12, FK12, GVB12, HM12, MSK^{+}12, O’B12, PNPP12, PH12, SH12a, Sei12, SSZ12, dLZ12]. EuroPKI [PNPP12]. Evaluation [Sch12]. EvoApplications [CAC^{+}12]. EvoBIO [GVB12]. EvoCOMNET [CAC^{+}12]. EvoCOMPLEX [CAC^{+}12]. EvoCOP [HM12]. EvoFIN [CAC^{+}12]. EvoGAMES [CAC^{+}12]. EvoHOT [CAC^{+}12]. EvoIASP [CAC^{+}12]. Evolutionary [CAC^{+}12, RKS^{+}12c, GVB12, HM12, MRC12]. EvoMUSART [MRC12]. EvoNUM [CAC^{+}12]. EvoPAR [CAC^{+}12]. EvoRISK [CAC^{+}12]. EvoSTIM [CAC^{+}12]. EvoSTOC [CAC^{+}12]. EWRL [SH12a]. EWSN [PH12]. Exact [MR12]. Experience [RSW12]. Experiments [JMP12].

Facing [KKW12]. FACS [BL12]. FASE [dLZ12]. Fault [Sch12]. FAW [SLSW12]. FAW-AAIM [SLSW12]. FC [DDS12, Dan12]. February [BLS12, DDS12, Dan12, DKN12, Dun12, HRB12, MDPQA12a, MDPQA12b, PH12, RiN12, RR12]. Ferrand [CP12]. Financial [DDS12, Dan12]. Finland [AJN12, RKM^{+}12]. First [DG12, HLKX12, KMMP12, MRC12, MOT12, ST12b].

FlashDB [YYH^{+}12]. FLOPS [ST12a]. FM [CP12]. FMCO [AdBB12]. FoIKS [LS12]. Fontainebleau [GPR^{+}12]. Formal [CP12, GTP12, GP12, MOT12, AdBB12, BL12]. Forum [OO12]. FOSSACS [Bir12]. Foundation [RD12]. Foundations [DKS12, Bir12, GAL12, LS12]. FPS [GAL12]. France [ADB^{+}12b, ADB^{+}12a, BJP12, BCC^{+}12, CP12, GAL12, GPR^{+}12, LV12, OO12, VMvN12]. Francisco [Dun12, KS12]. Frontiers [CHB^{+}12, SLSW12]. FSEN [AS12]. Functional [ST12a]. Fundamental [dLZ12]. Fundamentals [AS12]. Future [ÁCD^{+}12].

Games [vEV12]. GD [vKS12]. GECON [VAR12]. Gedi [BtHRBB12]. Genetic [MSK^{+}12]. Gent [VW12]. Geographical [MPT12]. Germany [ENO12, FBM12, HRB12, LS12, MR12, MP12, MK12, RD12, SH12b, Sch12, ST12b, SOP12]. GI [Sch12]. GI/ITG [Sch12]. Girona [AK12]. GPC [LCB12, RKM^{+}12]. Gran [MDPQA12a, MDPQA12b]. Graph [vKS12]. Graz [AdBB12]. Greece [GCFA12, KMMT12, SH12a]. Grid [BSW12, LCB12, RKM^{+}12]. Grids [VAR12]. Gros [Dan12]. Guimarães [BL12]. Gwangju [Ho12b, Ho12a].

Hagenberg [HNP12]. HAIS [CSA^{+}12b, CSA^{+}12a]. Hangzhou [RSW12]. Health [RtTM12, HLKX12]. Health-Care [RtTM12]. Heart [CKP^{+}12]. Held [BBK12, Bir12, CKP^{+}12, DG12, FK12, HK12, O’B12, RKM^{+}12, RKS^{+}12c, Sei12, VKG12, YSL12, dLZ12]. Heraklion [GCFA12]. HeteroPar [ADB^{+}12b, ADB^{+}12a]. HetsNets [BBK12]. HiBB [ADB^{+}12b, ADB^{+}12a]. Hiding [Shi12]. HIGHT [CWP12]. Hoc [ENO12]. Hong [BDDM12, CCAS12, LCB12]. Honor [ACÇ^{+}12]. HPCVirt [ADB^{+}12b, ADB^{+}12a]. HPPC [ADB^{+}12b, ADB^{+}12a]. HPSS [ADB^{+}12b, ADB^{+}12a]. HWTS [RKM^{+}12]. Hybrid [CSA^{+}12b, CSA^{+}12a].

IAPR [ST12b]. ICA [TCYZ12]. ICAISC [RKS^{+}12a, RKS^{+}12b, RKS^{+}12c]. ICDCIT [RR12]. ICDCN [BDDM12]. ICDEM [KA12]. Iceland [Jón12a, Jón12b]. ICIC [HGBF12, HGGG12, HGPH12]. ICWE [HK12]. IDP [WZH^{+}12]. IEKB [WZH^{+}12]. IFIP [BBK12, BKL^{+}12b, BKL^{+}12a, CK12, KH12, SSZ12]. II [KKW12, ADB^{+}12b, BKL^{+}12b, CSA^{+}12b, Gel12b, Ho12b, Jón12b, gLPZ^{+}12b, MDPQA12b, PCN12b, RKS^{+}12b, TCHB12a]. III [PCN12c]. Image [Ho12b, Ho12a]. Imaging [CKP^{+}12, YSL12]. Impossible [CWP12]. India [DLW12, Gel12a, Gel12b, KA12, KMMP12, RR12, TPCB12]. Indo [KMMP12]. Indo-Japan [KMMP12]. iNetSec [CK12]. Informatics [CVV12, FB12]. Information [BYdVZ^{+}12, BKL^{+}12c, HLKX12, LSW12, LS12, MPT12, PCN12c, PCN12a, PCN12b, PRK12, SLSW12, WJL^{+}12, AJN12, JY12, Lau12, LST^{+}12, RSW12]. Infrastructure [BSW12, LSW12]. Infrastructures [PNPP12]. Inspired [GTP12, MRC12, HGPH12]. Integrating [OO12]. Integration [BJP12]. Intelligence [BV12, HGGG12, KMMP12, RKS^{+}12a, RKS^{+}12b]. Intelligent [BKL^{+}12c, CSA^{+}12b, CSA^{+}12a, Gel12a, Gel12b, HGBF12, HGGG12, HGPH12, PCN12c, PCN12a, PCN12b]. Interaction [CBS^{+}12]. Interation [CBS^{+}12]. International [ABH12, AdBB12, AS12, BL12, BLS12, BBK12, BJP12, BKL^{+}12b, BKL^{+}12a, Bir12, BV12, BMSW12, BCC^{+}12, BDDM12, BKL^{+}12c, BtHRBB12, CKP^{+}12, CK12, CBS^{+}12, CHB^{+}12, CP12, CVTP12, Cho12, CCAS12, CVV12, CDN12, CSA^{+}12b, CSA^{+}12a, Dan12, DMV12, DG12, DLW12, DKN12, ENO12, FBM12, FK12, GANACBdCdV12, Gel12a, Gel12b, GPR^{+}12, GP12, HLKX12, HRB12, HNP12, HGBF12, HGGG12, HGPH12, IS12, Jón12a, Jón12b, JMP12, JP12, JY12, KA12, KS12, KBC^{+}12, KMMT12, KH12, KR12, LV12, gLPZ^{+}12b, gLPZ^{+}12a, LCB12, LST^{+}12, LS12, MRC12, MPT12, MR12, MFG12, MN12, MV12, MOT12, MDPQA12a, MDPQA12b, MK12, MGSM12, O’B12, OO12, PRK12, PSD12, PJ12, RiN12, RR12, RKM^{+}12, RD12, RZ12, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, RSW12, SSVW12, SH12b, Sch12, SMH^{+}12, ST12a, SA12]. International [SLSW12, SOP12, TR12, TCYZ12, TPCB12, VAR12, VMvN12, VSMS12, VMB^{+}12, VW12, VKG12, WJL^{+}12, WZH^{+}12, YGE12, YSL12, YYH^{+}12, dLZ12, vKS12]. Internet [ÁCD^{+}12, KMMT12, RR12]. Interpretation [KR12]. Invited [DKN12, SSVW12]. IPCAT [LST^{+}12]. IPEC [MR12]. IPM [AS12]. IR [BYdVZ^{+}12]. Iran [AS12]. ISBRA [BMSW12]. ISCA [VMvN12]. Island [JY12, MN12]. Islet [Dan12]. ISO605 [HS12]. ISPEC [RSW12]. Israel [BtHRBB12, TCYZ12]. Issue [GTP12]. Italy [Cra12, MPT12, MN12, PH12]. ITEMS [YYH^{+}12]. ITG [Sch12]. ITMAS [DHtM^{+}12]. IWSOS [KH12]. IX [SSVW12].

J [ACÇ^{+}12]. Jacques [Nac12]. January [BFG^{+}12, BDDM12, JMP12, KR12, KMMP12, RZ12, SMH^{+}12]. Japan [KMMP12, ST12a]. JBI [FN12]. Jean [Nac12]. Jean-Jacques [Nac12]. Jeju [JY12]. Joint [Bir12, BKL^{+}12c, DG12, FK12, MP12, O’B12, Sei12, SLSW12, dLZ12]. July [ABH12, AK12, CVV12, HNP12, KA12, Mac12, MOT12, MK12, OO12, RtTM12, SA12, VW12]. June [BCC^{+}12, BKL^{+}12c, BtHRBB12, CK12, CVV12, HK12, Jón12a, Jón12b, JP12, KMMT12, Mac12, MFG12, TCHB12b, TCHB12a, VMvN12]. June1 [BJP12].

Kaiserslautern [Sch12]. Kaohsiung [PCN12c, PCN12a, PCN12b]. Key [FBM12, PNPP12]. Kiel [LS12]. Klagenfurt [SMH^{+}12]. Knowledge [HKW12, LS12, TCHB12b, TCHB12a, RtTM12]. Knowledge-Centered [HKW12]. Kobe [ST12a]. Kolkata [DLW12, KMMP12]. Kong [BDDM12, CCAS12, LCB12]. Korea [Ho12b, Ho12a, JY12, gLPZ^{+}12b, gLPZ^{+}12a, YYH^{+}12]. Kozen [CS12]. KR4HC [RtTM12]. Krakow [SSZ12]. Kuala [TCHB12b, TCHB12a]. Kunming [SWJX12, WZH^{+}12].

Language [CVTP12, SA12, DMV12].

Languages [RZ12, SSVW12, Sei12]. Large [HKW12]. Large-Scale [HKW12]. LASER [MN12]. LATA [DMV12]. Latent [TCYZ12]. LATIN [FB12]. Learning [GVB12, SH12a, ST12b, VKG12]. Lectures [MN12]. Lednice [KBC^{+}12]. Lesná [ABH12]. Leuven [GANACBdCdV12, PNPP12]. Lightweight [CWP12]. Linear [BW12]. Linguistics [Gel12a, Gel12b]. Lithuania [VMB^{+}12]. Logic [ST12a, ACÇ^{+}12, BV12, CS12]. LPAR [BV12]. LPAR-18 [BV12]. Lucerne [CK12]. Lucia [DDS12, Dan12]. Lumpur [TCHB12b, TCHB12a]. LVA [TCYZ12]. LVA/ICA [TCYZ12].

MABS [VSMS12]. Machine [GVB12, KMMP12]. Málaga [GVB12, HM12, MRC12, MSK^{+}12, CAC^{+}12]. Malaysia [TCHB12b, TCHB12a]. Malo [VMvN12]. Management [GANACBdCdV12, KA12, MFG12, SLSW12, WJL^{+}12]. March [Bir12, BV12, CCAS12, CSA^{+}12b, CSA^{+}12a, Cra12, DDS12, Dan12, DMV12, DG12, Dun12, FK12, Gel12a, Gel12b, HRB12, KH12, LST^{+}12, LS12, MP12, O’B12, PCN12c, PCN12a, PCN12b, PSD12, RD12, Sch12, Sei12, TR12, TCYZ12, dLZ12]. Marek [ACÇ^{+}12]. Mathematical [ABH12, KBC^{+}12]. May [ACL12, BBK12, BJP12, BKL^{+}12b, BKL^{+}12a, BMSW12, BtHRBB12, CBS^{+}12, CHB^{+}12, CDN12, DHtM^{+}12, FBM12, GAL12, GCFA12, LCB12, PRK12, RKM^{+}12, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, SSVW12, SH12b, ST12a, SLSW12, TCHB12b, TCHB12a, VSMS12, VKG12]. MBC [WZH^{+}12]. MCBR [MGSM12]. MCBR-CDS [MGSM12]. MD [YGE12]. MDGS [ADB^{+}12b, ADB^{+}12a]. Measurement [TR12, Sch12]. Medical [MGSM12]. Medicine [FN12]. Membrane [GPR^{+}12]. MEMICS [KBC^{+}12]. Memorial [CVV12]. Mérida [BV12]. Methods [AdBB12, BtHRBB12, CP12, GTP12, GP12, KBC^{+}12]. MICCAI [CKP^{+}12, MGSM12, YSL12]. Mining [CBS^{+}12, CHB^{+}12, GVB12, TCHB12b, TCHB12a]. MITACS [GAL12]. Mlýn [BFG^{+}12]. MMB [Sch12]. MMCP [ABH12]. MMM [SMH^{+}12]. Mobile [ENO12]. Model [KR12]. Modeling [ABH12, AK12, OO12, SMH^{+}12, YGE12]. Modelling [CKP^{+}12, DKS12, Sch12]. MODELS [Kie12, ACL12, CKP^{+}12, JDK12, LSW12, Kie12]. Molecular [Cho12]. Monitoring [PSD12]. Multi [CDN12, DLW12, VSMS12]. Multi-Agent [CDN12, DLW12]. Multi-Agent-Based [VSMS12]. Multicore [KKW12]. Multidisciplinary [vEV12]. Multimedia [SMH^{+}12, Shi12]. Munich [HRB12]. Music [MRC12].

Nantes [BJP12]. Naples [MPT12]. NASA [GP12]. National [BSW12, Mac12]. Nature [GTP12]. Nature-Inspired [GTP12]. Netherlands [BLS12, KH12, vKS12]. Nets [JDK12]. Nets4Cars [VMB^{+}12]. Nets4Cars/Nets4Trains [VMB^{+}12]. Nets4Trains [VMB^{+}12]. Network [CK12]. NETWORKING [BBK12, BKL^{+}12b, BKL^{+}12a, BDDM12, TPCB12, BBK12, BKL^{+}12b, BKL^{+}12a]. Networks [ENO12, PH12]. NFM [GP12]. Nordic [AJN12, Lau12]. NordSec [AJN12, Lau12]. Norfolk [GP12]. Norms [ACÇ^{+}12]. November [AdBB12, DLW12, Ho12b, Ho12a, LV12]. Novosibirsk [CVV12]. Numeric [HNP12]. NV [PRK12].

Objects [AdBB12]. Occasion

[ACÇ^{+}12, CS12, DKN12, DKS12, Nac12]. October [AJN12, BL12, FN12, Kie12, KBC^{+}12, Lau12, SSZ12]. Online [SOP12]. Open [CK12]. Optimization [BJP12, HM12]. Organizing [KH12]. Other [JDK12]. Oulu [RKM^{+}12].

PA [JMP12, KR12, RZ12]. Pacific [Ho12b, Ho12a, SWJX12, TCHB12b, TCHB12a]. Pacific-Asia [TCHB12b, TCHB12a]. PADL [RZ12]. PAKDD [CHB^{+}12, TCHB12b, TCHB12a]. Palmas [MDPQA12a, MDPQA12b]. PAM [TR12]. Papers [ADB^{+}12b, ADB^{+}12a, Jón12a, Jón12b, MDPQA12a, MDPQA12b, ABH12, AdBB12, AS12, AK12, AJN12, BL12, BCC^{+}12, BKL^{+}12c, BtHRBB12, CKP^{+}12, CK12, CBS^{+}12, CHB^{+}12, CP12, CVV12, CDN12, DDS12, Dan12, DHtM^{+}12, DLW12, DKN12, ENO12, FN12, GANACBdCdV12, GAL12, GCFA12, GPR^{+}12, HK12, HNP12, HGBF12, HGGG12, HGPH12, IS12, JP12, JY12, KA12, KS12, Kie12, KBC^{+}12, LV12, Lau12, Mac12, MR12, MFG12, MV12, MP12, MOT12, MK12, MGSM12, OO12, PNPP12, RKM^{+}12, RtTM12, SSVW12, SH12a, ST12b, SA12, SOP12, SSZ12, TPCB12, VAR12, VMvN12, VSMS12, VW12, VKG12, WJL^{+}12, YSL12, vKS12].

Paphos [HK12, VAR12]. PARA [Jón12a, Jón12b]. Paradigms [KKW12]. Parallel [ADB^{+}12b, ADB^{+}12a, Jón12a, Jón12b, KKW12]. Parameterized [MR12]. Paris [GAL12, LV12]. Park [YGE12]. Part [ADB^{+}12b, ADB^{+}12a, BKL^{+}12b, BKL^{+}12a, Bir12, CSA^{+}12b, CSA^{+}12a, DG12, FK12, Gel12a, Gel12b, Ho12b, Ho12a, Jón12a, Jón12b, gLPZ^{+}12b, gLPZ^{+}12a, MDPQA12a, MDPQA12b, O’B12, PCN12c, PCN12a, PCN12b, RKS^{+}12a, RKS^{+}12b, Sei12, TCHB12b, TCHB12a, dLZ12]. Partially [ST12b]. Passive [TR12]. Perception [KMMP12]. PerMIn [KMMP12]. Personalized [FN12]. Perspectives [CVV12]. Peru [FB12]. Pervasive [LCB12, RKM^{+}12]. Petri [JDK12]. Philadelphia [JMP12, KR12, RZ12]. Physics [DKN12]. PKC [FBM12]. PL [BSW12]. Poland [BKL^{+}12c, RKS^{+}12a, RKS^{+}12b, RKS^{+}12c, SSZ12]. Portugal [BL12, CVTP12, SA12]. Portuguese [CVTP12]. POST [DG12]. Practical [RZ12]. Practice [BFG^{+}12, Bir12, DG12, DLW12, FBM12, FK12, GAL12, O’B12, PRK12, RSW12, Sei12, dLZ12]. Prague [BBK12, BKL^{+}12b, BKL^{+}12a]. Prediction [YGE12]. PRIMA [DLW12]. Principles [DG12, DLW12]. Privacy [GANACBdCdV12, JP12]. Problems [BJP12, CK12]. Proceedings [BKL^{+}12b, BKL^{+}12a, CSA^{+}12b, CSA^{+}12a, Gel12a, Gel12b, Ho12b, Ho12a, gLPZ^{+}12b, gLPZ^{+}12a, PCN12c, PCN12a, PCN12b, RKS^{+}12a, RKS^{+}12b, TCHB12b, TCHB12a, ACL12, BYdVZ^{+}12, BLS12, BBK12, BJP12, BFG^{+}12, Bir12, BV12, BMSW12, BDDM12, CVTP12, CAC^{+}12, Cho12, CCAS12, Cra12, DMV12, DG12, Dun12, FB12, FBM12, FK12, GVB12, GP12, HM12, HLKX12, HRB12, JMP12, KMMT12, KH12, KR12, KMMP12, LCB12, LST^{+}12, LS12, MRC12, MPT12, MSK^{+}12, O’B12, PRK12, PSD12, PH12, PJ12, RiN12, RR12, RD12, RZ12, RKS^{+}12c, RSW12, SH12b, Sch12, SMH^{+}12, ST12a, Sei12, SWJX12, SLSW12, TR12, TCYZ12, VMB^{+}12, WZH^{+}12, YGE12, YYH^{+}12, dLZ12]. Processing [ADB^{+}12b, ADB^{+}12a, CVTP12, Gel12a, Gel12b, LST^{+}12]. Program [CS12]. Programming [BJP12, BV12, MSK^{+}12, ST12a, Sei12, CDN12, Sei12]. Programs [ACÇ^{+}12]. ProMAS [CDN12]. Promises [ÁCD^{+}12]. ProPer [ADB^{+}12b, ADB^{+}12a]. PROPOR [CVTP12]. Protection [LSW12]. PSI [CVV12]. PSIVT [Ho12b, Ho12a]. PSL [ST12b]. Public [FBM12, FBM12, PNPP12].

Quality [RD12]. Quisquater [Nac12].

RC [VW12]. Reality [ÁCD^{+}12].

Reasoning [BV12]. Recognition [IS12]. RECOMB [Cho12]. Reconfigurable [CCAS12]. RED [LV12]. Reduced [BW12]. REFSQ [RD12]. Reinforcement [SH12a]. Reports [Kie12]. Representation [RtTM12]. Republic [BBK12, BKL^{+}12b, BKL^{+}12a, BFG^{+}12, KBC^{+}12]. Requirements [RD12]. Research [BYdVZ^{+}12, BMSW12, PRK12, Cho12]. Resilience [ADB^{+}12b, ADB^{+}12a]. Resource [LV12]. Results [HS12]. Retrieval [BYdVZ^{+}12, MGSM12]. Reversible [VW12]. Revised [ABH12, AdBB12, ADB^{+}12b, ADB^{+}12a, AS12, AK12, AJN12, BL12, BCC^{+}12, BKL^{+}12c, BtHRBB12, CKP^{+}12, CK12, CBS^{+}12, CHB^{+}12, CP12, CVV12, CDN12, DDS12, Dan12, DHtM^{+}12, DLW12, DKN12, ENO12, FN12, GANACBdCdV12, GAL12, GCFA12, GPR^{+}12, HK12, HNP12, HGBF12, HGGG12, HGPH12, IS12, Jón12a, Jón12b, JP12, JY12, KA12, KS12, Kie12, KBC^{+}12, LV12, Lau12, Mac12, MR12, MFG12, MN12, MV12, MP12, MOT12, MDPQA12a, MDPQA12b, MK12, MGSM12, OO12, PNPP12, RKM^{+}12, RtTM12, SSVW12, SH12a, ST12b, SA12, SOP12, SSZ12, TPCB12, VAR12, VMvN12, VSMS12, VW12, VKG12, WJL^{+}12, YSL12, vKS12]. Reykjavík [Jón12a, Jón12b]. RFID. [JP12]. RFIDSec [JP12]. Rim [Ho12b, Ho12a]. RLCPS [DDS12]. Rodney [DDS12]. RSA [Dun12]. Runtime [KS12]. Russia [CVV12]. RV [KS12].

S [DKN12]. S3E [RKM^{+}12]. Saarbrücken [ENO12, MR12, MP12, SOP12]. SAC [MV12]. Saint [VMvN12]. Saint-Malo [VMvN12]. Salamanca [CSA^{+}12b, CSA^{+}12a]. San [Dun12, KS12]. Santorini [KMMT12]. SBP [YGE12]. Scale [HKW12, BtHRBB12]. Schools [MN12]. Science [ABH12, BFG^{+}12, Bir12, DKN12, GTP12, HLKX12, KBC^{+}12, PRK12]. Scientific [BSW12, Jón12a, Jón12b]. SDL [OO12, OO12]. Second [CKP^{+}12, KA12, KS12, MGSM12]. Secure [AJN12, BLS12, Lau12, SH12b]. Security [AJN12, CK12, DDS12, Dan12, DG12, GANACBdCdV12, GAL12, JP12, JY12, Lau12, Mac12, MP12, Nac12, RSW12, Shi12, TPCB12, BKL^{+}12c, MFG12]. Selected [ABH12, ADB^{+}12b, ADB^{+}12a, AS12, AK12, AJN12, BL12, BCC^{+}12, BKL^{+}12c, BtHRBB12, CKP^{+}12, CK12, CBS^{+}12, CHB^{+}12, CP12, CVV12, CDN12, DDS12, Dan12, DHtM^{+}12, DLW12, DKN12, ENO12, FN12, GANACBdCdV12, GAL12, GCFA12, GPR^{+}12, HS12, HK12, HNP12, HGBF12, HGGG12, HGPH12, IS12, Jón12a, Jón12b, JP12, JY12, KA12, KS12, Kie12, KBC^{+}12, LV12, Lau12, Mac12, MR12, MFG12, MV12, MP12, MOT12, MDPQA12a, MDPQA12b, MK12, MGSM12, PNPP12, RKM^{+}12, RtTM12, SSVW12, SH12a, ST12b, SA12, SOP12, SSZ12, TPCB12, VAR12, VMvN12, VSMS12, VKG12, WJL^{+}12, YSL12, vKS12, MV12]. Self [KH12]. Self-Organizing [KH12]. Semantic [GCFA12]. Semantics [CS12]. SenDe [WZH^{+}12]. Sensor [ENO12, PH12]. Separation [TCYZ12]. September [ADB^{+}12b, ADB^{+}12a, CKP^{+}12, CP12, ENO12, GANACBdCdV12, IS12, KS12, MR12, MGSM12, PNPP12, SH12a, ST12b, SOP12, WJL^{+}12, YSL12, vKS12]. Sergot [ACÇ^{+}12]. Services [CP12, PNPP12, VAR12]. SET [SSZ12]. SETOP [GANACBdCdV12]. Shenzhen [CHB^{+}12]. Sicily [Cra12]. SIDE [RKS^{+}12c, SH12b]. Side-Channel [SH12b]. Signal [TCYZ12]. SIIS [BKL^{+}12c]. SIM [YYH^{+}12]. Simulation [VSMS12]. SLE [SA12]. Slovakia [ABH12]. Slovenia [RtTM12]. SNA [WJL^{+}12]. SNSM [YYH^{+}12]. Social [vEV12, YGE12]. SOFSEM [BFG^{+}12]. Soft [RKS^{+}12a, RKS^{+}12b]. Software [AS12, BL12, BLS12, Bir12, DG12, FK12, JMP12, Kie12, MN12, O’B12, OO12, RD12, Sei12, SSZ12, dLZ12, vEV12, SA12]. Sound [MRC12]. South [Ho12b, Ho12a, gLPZ^{+}12b, gLPZ^{+}12a, YYH^{+}12]. Space [BtHRBB12]. Spain [AK12, BYdVZ^{+}12, CAC^{+}12, Cho12, CSA^{+}12b, CSA^{+}12a, DMV12, FN12, GVB12, HM12, MRC12, MOT12, MSK^{+}12, MDPQA12a, MDPQA12b]. Spanish [FN12]. Special [GTP12]. Spindleruv [BFG^{+}12]. Spontaneous [GANACBdCdV12]. SSVM [BtHRBB12]. St [DDS12, Dan12]. STACOM [CKP^{+}12]. Stará [ABH12]. Statistical [CKP^{+}12]. STM [MFG12]. Structures [Bir12]. Summer [MN12]. Supervised [ST12b]. Support [MGSM12]. Surathkal [TPCB12]. Surfaces [BCC^{+}12]. Swarm [RKS^{+}12c]. Switzerland [CK12]. Symposia [Kie12, RKS^{+}12c]. Symposium [AdBB12, BLS12, BMSW12, CCAS12, ENO12, FB12, FN12, GP12, HK12, Ho12b, Ho12a, LS12, MPT12, MR12, RZ12, ST12a, Sei12, vKS12]. System [OO12]. Systems [AJN12, BLS12, BKL^{+}12c, CVV12, CDN12, CSA^{+}12b, CSA^{+}12a, DLW12, ENO12, FK12, HKW12, HRB12, KH12, Lau12, gLPZ^{+}12b, gLPZ^{+}12a, LS12, MPT12, MDPQA12a, MDPQA12b, PCN12c, PCN12a, PCN12b, PRK12, Sch12, Sei12, VAR12, YYH^{+}12, GTP12].

TACAS [FK12]. TAFA [MOT12]. Taipei [CBS^{+}12, DHtM^{+}12, SSVW12, VSMS12, VKG12]. Taiwan [CBS^{+}12, DHtM^{+}12, PCN12c, PCN12a, PCN12b, SSVW12, VSMS12, VKG12]. Tallinn [Bir12, DG12, FK12, Lau12, O’B12, Sei12, dLZ12]. TAMC [ACL12]. Taormina [Cra12]. TC [BBK12, BKL^{+}12b, BKL^{+}12a, KH12, SSZ12]. TC3 [ST12b]. TCC [Cra12]. TEA [CWP12]. Technical [BSW12]. Techniques [BJP12, MK12, PJ12, SSZ12]. Technologies [KKW12, SSVW12, SWJX12, VMB^{+}12, WZH^{+}12]. Technology [AJN12, DHtM^{+}12, Ho12b, Ho12a, Lau12, RR12]. Tehran [AS12]. Tel [TCYZ12, HS12]. Telecommunication [HS12]. Text [Gel12a, Gel12b]. Thalheim [DKS12]. Theoretical [DKN12, DKS12, FB12]. Theorie [MOT12]. Theories [HGGG12, JMP12]. Theory [BFG^{+}12, Bir12, Cra12, DMV12, DG12, FBM12, FK12, MDPQA12a, MDPQA12b, Nac12, O’B12, PRK12, PJ12, Sei12, dLZ12, ACL12, Cra12, MP12]. Third [BtHRBB12, LV12, SH12b, VW12, YSL12]. Tiruchirappalli [KA12]. Tissues [LST^{+}12]. TMA [PSD12]. Tolerance [Sch12]. Tools [CCAS12, JMP12, FK12]. Topics [Dun12]. Toronto [CKP^{+}12, CDN12, MV12, MGSM12, YSL12]. Torremolinos [FN12]. TOSCA [MP12]. Toulouse [OO12]. Track [Dun12]. Traffic [PSD12]. Transactions [GTP12, HKW12, JDK12, PCM^{+}12, Shi12]. Trends [BFG^{+}12, HK12, MK12]. Trento [PH12]. Trust [DG12, MFG12]. Tutorial [MN12]. Tutorials [HK12]. TX [BMSW12].

UCHPC [ADB^{+}12b, ADB^{+}12a]. UK [LST^{+}12, Mac12, PJ12]. Ulm [ST12b]. UMAP [AK12]. USA [BMSW12, Dun12, GP12, JMP12, JP12, KS12, KR12, PRK12, RZ12, YGE12]. User [AK12].

V [HKW12, JDK12]. VA [GP12]. Variable [TCYZ12]. Variational [BtHRBB12]. Vegas [PRK12]. Vehicles [VMB^{+}12]. Venezuela [BV12]. Verification [KS12, MN12, KR12]. Verified [JMP12]. VHPC [ADB^{+}12b, ADB^{+}12a]. Video [Ho12b, Ho12a]. Vienna [PSD12, TR12]. VII [PCM^{+}12, Shi12]. Vilnius [VMB^{+}12]. Vision [BtHRBB12]. VMCAI [KR12]. VSTTE [JMP12].

W2GIS [MPT12]. WADT [MK12]. WAIM [WJL^{+}12]. WALCOM [RiN12, RiN12]. WAOA [SOP12]. Warsaw [BKL^{+}12c]. Web [GCFA12, CP12, HK12, MPT12, SWJX12, WJL^{+}12, WZH^{+}12]. Web-Age [WJL^{+}12]. WECSR [DDS12]. WEED [VMvN12]. Wellington [Kie12]. WG [CK12]. WGIM [WJL^{+}12]. WIOSCA [VMvN12]. Wired [KMMT12]. Wired/Wireless [KMMT12]. Wireless [ENO12, MPT12, KMMT12, PH12]. WISA [JY12]. Working [RD12]. Workshop [BL12, CKP^{+}12, CK12, CBS^{+}12, CP12, CDN12, DKN12, GANACBdCdV12, GAL12, IS12, JP12, JY12, KBC^{+}12, KH12, LV12, MFG12, MV12, MP12, MOT12, MK12, MGSM12, PSD12, PNPP12, RiN12, RtTM12, SSVW12, SH12a, SH12b, ST12b, SOP12, VAR12, VSMS12, VMB^{+}12, VW12, VKG12, YSL12]. Workshops [ADB^{+}12b, ADB^{+}12a, AK12, BBK12, CHB^{+}12, DDS12, DHtM^{+}12, GCFA12, HK12, Kie12, RKM^{+}12, VMvN12, WJL^{+}12, WZH^{+}12, YYH^{+}12]. WS [CP12]. WS-FM [CP12]. WTCS [DKN12]. Wuhan [WJL^{+}12]. WWIC [KMMT12].

XII [VSMS12]. XMLDM [WJL^{+}12]. XTEA [CWP12]. XV [GTP12].

Zakopane [RKS^{+}12a, RKS^{+}12b, RKS^{+}12c]. Zealand [DKN12, Kie12]. Zero [BW12]. Zero-Correlation [BW12]. Zhengzhou [HGBF12, HGGG12, HGPH12].

References

Adam:2012:MMC

[ABH12] Gheorghe Adam, Ján Busa, and Michal Hnatic, editors. Mathematical Modeling and Computational Science: International Conference, MMCP 2011, Stará Lesná, Slovakia, July 4– 8, Revised Selected Papers, volume 7125 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28211-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28211-9.

Artikis:2012:LPN

[ACÇ^{+}12] Alexander Artikis, Robert Craven, Nihan Kesim Çiçekli, Babak Sadighi, and Kostas Stathis, editors. Logic Programs, Norms and Action: Essays in Honor of Marek J. Sergot on the Occasion of His 60th Birthday, volume 7360 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29413-8. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-29413-6.

Alvarez:2012:FIF

[ÁCD^{+}12] Federico Álvarez, Frances Cleary, Petros Daras, John Domingue, Alex Galis, and et al, editors. The Future Internet: Future Internet Assembly 2012: From Promises to Reality, volume 7281 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30240-8. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-30240-4.

Agrawal:2012:TAM

[ACL12] Manindra Agrawal, S. Barry Cooper, and Angsheng Li, editors. Theory and Applications of Models of Computation: 9th Annual Conference, TAMC 2012, Beijing, China, May 16–21. Proceedings, volume 7287 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29951-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29951-3.

Alexander:2012:EPPb

[ADB^{+}12a] Michael Alexander, Pasqua D’Ambra, Adam Belloum, George Bosilca, Mario Cannataro, and et al, editors. Euro-Par 2011: Parallel Processing Workshops: CCPI, CGWS, HeteroPar, HiBB, HPCVirt, HPPC, HPSS, MDGS, ProPer, Resilience, UCHPC, VHPC, Bordeaux, France, August 29 — September 2, 2011, Revised Selected Papers, Part I, volume 7155 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29736-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29736-6.

Alexander:2012:EPPa

[ADB^{+}12b] Michael Alexander, Pasqua D’Ambra, Adam Belloum, George Bosilca, Mario Cannataro, and et al, editors. Euro-Par 2011: Parallel Processing Workshops: CCPI, CGWS, HeteroPar, HiBB, HPCVirt, HPPC, HPSS, MDGS, ProPer, Resilience, UCHPC, VHPC, Bordeaux, France, August 29 — September 2, 2011, Revised Selected Papers, Part II, volume 7156 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29739-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29739-7.

Aichernig:2012:FMC

[AdBB12] Bernhard K. Aichernig, Frank S. de Boer, and Marcello M. Bonsangue, editors. Formal Methods for Components and Objects: 9th International Symposium, FMCO 2010, Graz, Austria, November 29 — December 1, 2010. Revised Papers, volume 6957 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25270-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25270-9.

Aura:2012:IST

[AJN12] Tuomas Aura, Kimmo Järvinen, and Kaisa Nyberg, editors. Information Security Technology for Applications: 15th Nordic Conference on Secure IT Systems, NordSec 2010, Espoo, Finland, October 27– 29, Revised Selected Papers, volume 7127 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27936-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27936-2.

Ardissono:2012:AUM

[AK12] Liliana Ardissono and Tsvi Kuflik, editors. Advances in User Modeling: UMAP 2011 Workshops, Girona, Spain, July 11– 15, Revised Selected Papers, volume 7138 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28508-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28508-0.

Arbab:2012:FSE

[AS12] Farhad Arbab and Marjan Sirjani, editors. Fundamentals of Software Engineering: 4th IPM International Conference, FSEN 2011, Tehran, Iran, April 20– 22, Revised Selected Papers, volume 7141 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29319-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29319-1.

Becvar:2012:NWI

[BBK12] Zdenek Becvar, Robert Bestak, and Lukas Kencl, editors. NETWORK-ING 2012 Workshops: International IFIP TC 6 Workshops, ETICS, HetsNets, and CompNets, Held at NETWORKING 2012, Prague, Czech Republic, May 25, 2012. Proceedings, volume 7291 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30038-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
30038-7.

Boissonnat:2012:CSI

[BCC^{+}12] Jean-Daniel Boissonnat, Patrick Chenin, Albert Cohen, Christian Gout, Tom Lyche, and et al, editors. Curves and Surfaces: 7th International Conference, Avignon, France, June 24 — 30, 2010, Revised Selected Papers, volume 6920 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27412-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27412-1.

Bononi:2012:DCN

[BDDM12] Luciano Bononi, Ajoy K.

Datta, Stéphane Devismes, and Archan Misra, editors. Distributed Computing and Networking: 13th International Conference, ICDCN 2012, Hong Kong, China, January 3–6. Proceedings, volume 7129 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25958-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25958-6.

Bielikova:2012:STP

[BFG^{+}12] Mária Bieliková, Gerhard Friedrich, Georg Gottlob, Stefan Katzenbeisser, and György Turán, editors. SOF-SEM 2012: Theory and Practice of Computer Science: 38th Conference on Current Trends in Theory and Practice of Computer Science, Spindleruv Mlýn, Czech Republic, January 21–27. Proceedings, volume 7147 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27659-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27659-0.

Birkedal:2012:FSS

[Bir12] Lars Birkedal, editor. Foundations of Software Science and Computational Structures: 15th International Conference, FOSSACS 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7213 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28728-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28728-2.

Beldiceanu:2012:IAT

[BJP12] Nicolas Beldiceanu, Narendra Jussien, and Éric Pinson, editors. Integration of AI and OR Techniques in Contraint Programming for Combinatorial Optimization Problems: 9th International Conference, CPAIOR 2012, Nantes, France, May 28 — June1, 2012. Proceedings, volume 7298 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29827-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        29827-1.

Bestak:2012:NIIb

[BKL^{+}12a] Robert Bestak, Lukas Kencl, Li Erran Li, Joerg Widmer, and Hao Yin, editors. NET-WORKING 2012: 11th International IFIP TC 6 Networking Conference, Prague, Czech Republic, May 21–25, Proceedings, Part I, volume 7289 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30044-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        30044-8.

Bestak:2012:NIIa

[BKL^{+}12b] Robert Bestak, Lukas Kencl, Li Erran Li, Joerg Widmer, and Hao Yin, editors. NET-WORKING 2012: 11th International IFIP TC 6 Networking Conference, Prague, Czech Republic, May 21–25, Proceedings, Part II, volume 7290 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30053-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
30053-0.

Bouvry:2012:SII

[BKL^{+}12c] Pascal Bouvry, Mieczyslaw A. Klopotek, Franck Leprévost, Malgorzata Marciniak, Agnieszka Mykowiecka, and et al, editors. Security and Intelligent Information Systems: International Joint Conferences, SIIS 2011, Warsaw, Poland, June 13–14, Revised Selected Papers, volume 7053 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25260-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25260-0.

Barbosa:2012:FAC

[BL12] Luís Soares Barbosa and Markus Lumpe, editors. Formal Aspects of Component Software: 7th International Workshop, FACS 2010, Guimarães, Portugal, October 14– 16, Revised Selected Papers, volume 6921 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27268-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27268-4.

Barthe:2012:ESS

[BLS12] Gilles Barthe, Benjamin Livshits, and Riccardo Scandariato, editors. Engineering Secure Software and Systems: 4th International Symposium, ESSoS 2012, Eindhoven, The Netherlands, February, 16–17. Proceedings, volume 7159 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28165-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28165-5.

Bleris:2012:BRA

[BMSW12] Leonidas Bleris, Ion Mandoiu, Russell Schwartz, and Jianxin Wang, editors. Bioinformatics Research and Applications: 8th International Symposium, IS-BRA 2012, Dallas, TX, USA, May 21–23. Proceedings, volume 7292 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30190-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
30190-2.

Bubak:2012:BND

[BSW12] Marian Bubak, Tomasz Szepieniec, and Kazimierz Wiatr, editors. Building a National Distributed e-Infrastructure-PL-Grid: Scientific and Technical Achievements, volume 7136 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28266-0. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-28266-9.

Bruckstein:2012:SSV

[BtHRBB12] Alfred M. Bruckstein, Bart M. ter Haar Romeny, Alexander M. Bronstein, and Michael M. Bronstein, editors. Scale Space and Variational Methods in Computer Vision: Third International Conference, SSVM 2011, Ein-Gedi, Israel, May 29 — June 2, 2011, Revised Selected Papers, volume 6667 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-24784-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24784-2.

Bjorner:2012:LPA

[BV12] Nikolaj Bjørner and Andrei Voronkov, editors. Logic for Programming, Artificial Intelligence, and Reasoning: 18th International Conference, LPAR-18, Mérida, Venezuela, March 11–15. Proceedings, volume 7180 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28716-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28716-9.

Bogdanov:2012:ZCL

[BW12] Andrey Bogdanov and Meiqin Wang. Zerocorrelation linear cryptanalysis with reduced data complexity. Lecture Notes in Computer Science, 7549:29–48, 2012.

Baeza-Yates:2012:AIR

[BYdVZ^{+}12] Ricardo Baeza-Yates, Arjen P. de Vries, Hugo Zaragoza, B. Barla Cambazoglu, Vanessa Murdock, and et al, editors. Advances in Information Retrieval: 34th European Conference on IR Research, ECIR 2012, Barcelona, Spain, April 1–5. Proceedings, volume 7224 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28996-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28996-5.

Chio:2012:AEC

[CAC^{+}12] Cecilia Di Chio, Alexandros Agapitos, Stefano Cagnoni, Carlos Cotta, Francisco Fernández de Vega, and et al, editors. Applications of Evolutionary Computation: EvoApplications 2012: EvoCOM-NET, EvoCOMPLEX, EvoFIN, EvoGAMES, EvoHOT, EvoIASP, EvoNUM, EvoPAR, EvoRISK, EvoS-TIM, and EvoSTOC, Málaga, Spain, April 11–13, Proceedings, volume 7248 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29177-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29177-7.

Cao:2012:ADM

[CBS^{+}12] Longbing Cao, Ana L. C. Bazzan, Andreas L. Symeonidis, Vladimir I. Gorodetsky, Gerhard Weiss, and et al, editors. Agents and Data Mining Interaction: 7th International Workshop on Agents and Data Mining Interation, ADMI 2011, Taipei, Taiwan, May 2– 6, Revised Selected Papers, volume 7103 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27608-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27608-8.

Choy:2012:RCA

[CCAS12] Oliver C. S. Choy, Ray C. C. Cheung, Peter Athanas, and Kentaro Sano, editors. Reconfigurable Computing: Architectures, Tools and Applications: 8th International Symposium, ARC 2012, Hong Kong, China, March 19–23. Proceedings, volume 7199 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28364-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28364-2.

Collier:2012:PMA

[CDN12] Rem Collier, Jürgen Dix, and Peter Novák, editors. Programming Multi-Agent Systems: 8th International Workshop, ProMAS 2010, Toronto, ON, Canada, May 11, 2010. Revised Selected Papers, volume 6599 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28938-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28938-5.

Cao:2012:NFA

[CHB^{+}12] Longbing Cao, Joshua Zhexue Huang, James Bailey, Yun Sing Koh, and Jun Luo, editors. New Frontiers in Applied Data Mining: PAKDD 2011 International Workshops, Shenzhen, China, May 24–27, Revised Selected Papers, volume 7104 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28319-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28319-2.

Chor:2012:RCM

[Cho12] Benny Chor, editor. Research in Computational Molecular Biology: 16th Annual International Conference, RE-COMB 2012, Barcelona, Spain, April 21–24. Proceedings, volume 7262 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29626-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29626-0.

Camenisch:2012:OPN

[CK12] Jan Camenisch and Dogan Kesdogan, editors.

Open Problems in Network Security: IFIP WG 11.4 International Workshop, iNetSec 2011, Lucerne, Switzerland, June 9, 2011, Revised Selected Papers, volume 7039 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27584-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27584-5.

Camara:2012:SAC

[CKP^{+}12] Oscar Camara, Ender Konukoglu, Mihaela Pop, Kawal Rhode, Maxime Sermesant, and et al, editors. Statistical Atlases and Computational Models of the Heart. Imaging and Modelling Challenges: Second International Workshop, STA-COM 2011, Held in Conjunction with MIC-CAI 2011, Toronto, ON, Canada, September 22, 2011, Revised Selected Papers, volume 7085 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28325-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28325-3.

Carbone:2012:WSF

[CP12] Marco Carbone and Jean-Marc Petit, editors. Web Services and Formal Methods: 8th International Workshop, WS-FM 2011, Clermont-Ferrand, France, September 1–2, Revised Selected Papers, volume 7176 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29833-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29833-2.

Cramer:2012:TCT

[Cra12] Ronald Cramer, editor. Theory of Cryptography: 9th Theory of Cryptography Conference, TCC 2012, Taormina, Sicily, Italy, March 19–21. Proceedings, volume 7194 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28913-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.

com/content/978-3-642-
28913-2.

Constable:2012:LPS

[CS12] Robert L. Constable and Alexandra Silva, editors. Logic and Program Semantics: Essays Dedicated to Dexter Kozen on the Occasion of His 60th Birthday, volume 7230 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29484-7. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-29484-6.

Corchado:2012:HAIb

[CSA^{+}12a] Emilio Corchado, Václav Snásel, Ajith Abraham, Michal Wozniak, Manuel Graña, and et al, editors. Hybrid Artificial Intelligent Systems: 7th International Conference, HAIS 2012, Salamanca, Spain, March 28-30th, 2012. Proceedings, Part I, volume 7208 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28941-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        28941-5.

Corchado:2012:HAIa

[CSA^{+}12b] Emilio Corchado, Václav Snásel, Ajith Abraham, Michal Wozniak, Manuel Graña, and et al, editors. Hybrid Artificial Intelligent Systems: 7th International Conference, HAIS 2012, Salamanca, Spain, March 28-30th, 2012. Proceedings, Part II, volume 7209 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28930-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28930-9.

Caseli:2012:CPP

[CVTP12] Helena Caseli, Aline Villavicencio, António Teixeira, and Fernando Perdigão, editors. Computational Processing of the Portuguese Language: 10th International Conference, PRO-POR 2012, Coimbra, Portugal, April 17–20. Proceedings, volume 7243 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28884-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28884-5.

Clarke:2012:PSI

[CVV12] Edmund Clarke, Irina Virbitskaite, and Andrei Voronkov, editors. Perspectives of Systems Informatics: 8th International Andrei Ershov Memorial Conference, PSI 2011, Novosibirsk, Russia, June 27–July 1, 2011, Revised Selected Papers, volume 7162 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29708-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29708-3.

Chen:2012:IDC

[CWP12] Jiazhe Chen, Meiqin Wang, and Bart Preneel. Impossible differential cryptanalysis of the lightweight block ciphers TEA, XTEA and HIGHT. Lecture Notes in Computer Science, 7374:117– 137, 2012. URL http:

//link.springer.com/

chapter/10.1007/978-
3-642-31410-0_8.

Danezis:2012:FCDb

[Dan12] George Danezis, editor. Financial Cryptography and Data Security: 15th International Conference, FC 2011, Gros Islet, St. Lucia, February 28 — March 4, 2011, Revised Selected Papers, volume 7035 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27575-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27575-3.

Danezis:2012:FCDa

[DDS12] George Danezis, Sven Dietrich, and Kazue Sako, editors. Financial Cryptography and Data Security: FC 2011 Workshops, RLCPS and WECSR 2011, Rodney Bay, St. Lucia, February 28 — March 4, 2011, Revised Selected Papers, volume 7126 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29888-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29888-2.

Degano:2012:PST

[DG12] Pierpaolo Degano and Joshua D. Guttman, editors. Principles of Security and Trust: First International Conference, POST 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012, Proceedings, volume 7215 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28640-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28640-7.

Dechesne:2012:AAT

[DHtM^{+}12] Francien Dechesne, Hiromitsu Hattori, Adriaan ter Mors, Jose Miguel Such, Danny Weyns, and et al, editors. Advanced Agent Technology: AA-MAS 2011 Workshops, AMPLE, AOSE, ARMS, DOCM^{3} AS, ITMAS, Taipei, Taiwan, May 2– 6. Revised Selected Papers, volume 7068 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27215-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27215-8.

Dinneen:2012:CPB

[DKN12] Michael J. Dinneen, Bakhadyr Khoussainov, and André Nies, editors. Computation, Physics and Beyond: International Workshop on Theoretical Computer Science, WTCS 2012, Dedicated to Cristian S. Calude on the Occasion of His 60th Birthday, Auckland, New Zealand, February 21–24, Revised Selected and Invited Papers, volume 7160 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27653-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27653-8.

Dusterhoft:2012:CMT

[DKS12] Antje Düsterhöft, Meike Klettke, and Klaus-Dieter Schewe, editors. Conceptual Modelling and Its Theoretical Foundations: Essays Dedicated to Bernhard Thalheim on the Occasion of His 60th Birthday, volume 7260 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28278-4. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-28278-2.

Desai:2012:PPM

[DLW12] Nirmit Desai, Alan Liu, and Michael Winikoff, editors. Principles and Practice of Multi-Agent Systems: 13th International Conference, PRIMA 2010, Kolkata, India, November 12– 15, Revised Selected Papers, volume 7057 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25919-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25919-7.

deLara:2012:FAS

[dLZ12] Juan de Lara and Andrea Zisman, editors. Fundamental Approaches to Software Engineering: 15th International Conference, FASE 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7212 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28871-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28871-5.

Dediu:2012:LAT

[DMV12] Adrian-Horia Dediu and Carlos Martín-Vide, editors. Language and Automata Theory and Applications: 6th International Conference, LATA 2012, A Coruña, Spain, March 5–9. Proceedings, volume 7183 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28331-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.

com/content/978-3-642-
28331-4.

Dunkelman:2012:TCC

[Dun12] Orr Dunkelman, editor. Topics in Cryptology — CT-RSA 2012: The Cryptographers’ Track at the RSA Conference 2012, San Francisco, CA, USA, February 27 — March 2, 2012. Proceedings, volume 7178 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27953-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27953-9.

Erlebach:2012:ASS

[ENO12] Thomas Erlebach, Sotiris Nikoletseas, and Pekka Orponen, editors. Algorithms for Sensor Systems: 7th International Symposium on Algorithms for Sensor Systems, Wireless Ad Hoc Networks and Autonomous Mobile Entities, ALGOSENSORS 2011, Saarbrücken, Germany, September 8–9, Revised Selected Papers, volume 7111 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28208-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28208-9.

Fernandez-Baca:2012:LTI

[FB12] David Fernández-Baca, editor. LATIN 2012: Theoretical Informatics: 10th Latin American Symposium, Arequipa, Peru, April 16–20. Proceedings, volume 7256 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29343-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29343-6.

Fischlin:2012:PKC

[FBM12] Marc Fischlin, Johannes Buchmann, and Mark Manulis, editors. Public Key Cryptography — PKC 2012: 15th International Conference on Practice and Theory in Public Key Cryptography, Darmstadt, Germany, May 21–23. Proceedings, volume 7293 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30056-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
30056-1.

Flanagan:2012:TAC

[FK12] Cormac Flanagan and Barbara König, editors. Tools and Algorithms for the Construction and Analysis of Systems: 18th International Conference, TACAS 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7214 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28755-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28755-8.

Freitas:2012:BPM

[FN12] Ana T. Freitas and Arcadi Navarro, editors. Bioinformatics for Personalized Medicine: 10th Spanish Symposium, JBI 2010, Torremolinos, Spain, October 27– 29. Revised Selected Papers, volume 6620 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28061-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28061-0.

Garcia-Alfaro:2012:FPS

[GAL12] Joaquin Garcia-Alfaro and Pascal Lafourcade, editors. Foundations and Practice of Security: 4th Canada-France MI-TACS Workshop, FPS 2011, Paris, France, May 12–13, Revised Selected Papers, volume 6888 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27900-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27900-3.

Garcia-Alfaro:2012:DPM

[GANACBdCdV12] Joaquin Garcia-Alfaro, Guillermo Navarro-Arribas, Nora Cuppens-Boulahia, and Sabrina de Capitani di Vimercati, editors. Data Privacy Management and Autonomous Spontaneous Security: 6th International Workshop, DPM 2011, and 4th International Workshop, SE-TOP 2011, Leuven, Belgium, September 15–16, Revised Selected Papers, volume 7122 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28878-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28878-4.

Garcia-Castro:2012:SWE

[GCFA12] Raúl García-Castro, Dieter Fensel, and Grigoris Antoniou, editors. The Semantic Web: ESWC 2011 Workshops: ESWC 2011 Workshops, Heraklion, Greece, May 29– 30, Revised Selected Papers, volume 7117 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25952-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25952-4.

Gelbukh:2012:CLIa

[Gel12a] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 13th International Conference, CICLing 2012, New Delhi, India, March 11–17, Proceedings, Part I, volume 7181 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28603-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28603-2.

Gelbukh:2012:CLIb

[Gel12b] Alexander Gelbukh, editor. Computational Linguistics and Intelligent Text Processing: 13th International Conference, CICLing 2012, New Delhi, India, March 11–17, Proceedings, Part II, volume 7182 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28600-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28600-1.

Lee:2012:DSAb

[gLPZ^{+}12a] Sang goo Lee, Zhiyong Peng, Xiaofang Zhou, Yang-Sae Moon, Rainer Unland, and et al, editors. Database Systems for Advanced Applications: 17th International Conference, DASFAA 2012, Busan, South Korea, April 15– 19, Proceedings, Part I, volume 7238 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29037-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29037-4.

Lee:2012:DSAa

[gLPZ^{+}12b] Sang goo Lee, Zhiyong Peng, Xiaofang Zhou, Yang-Sae Moon, Rainer Unland, and et al, editors. Database Systems for Advanced Applications: 17th International Conference, DASFAA 2012, Busan, South Korea, April 15– 19, Proceedings, Part II, volume 7239 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29034-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29034-3.

Goodloe:2012:NFM

[GP12] Alwyn E. Goodloe and Suzette Person, editors. NASA Formal Methods: 4th International Symposium, NFM 2012, Norfolk, VA, USA, April 3–5. Proceedings, volume 7226 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28890-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28890-6.

Gheorghe:2012:MCI

[GPR^{+}12] Marian Gheorghe, Gheorghe Paun, Grzegorz Rozenberg, Arto Salomaa, and Sergey Verlan, editors. Membrane Computing: 12th International Conference, CMC 2011, Fontainebleau, France, August 23–26, Revised Selected Papers, volume 7184 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28023-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28023-8.

Gavrilova:2012:TCS

[GTP12] Marina L. Gavrilova, C. J. Kenneth Tan, and Cong-Vinh Phan, editors. Transactions on Computational Science XV: Special Issue on Advances in Autonomic Computing: Formal Engineering Methods for Nature-Inspired Computing Systems, volume 7050 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28524-4. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-28524-0.

Giacobini:2012:ECM

[GVB12] Mario Giacobini, Leonardo Vanneschi, and William S. Bush, editors. Evolutionary Computation, Machine Learning and Data Mining in Bioinformatics: 10th European Conference, EvoBIO 2012, Málaga, Spain, April 11–13. Proceedings, volume 7246 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29065-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29065-7.

Huang:2012:AICa

[HGBF12] De-Shuang Huang, Yong Gan, Vitoantonio Bevilacqua, and Juan Carlos Figueroa, editors. Advanced Intelligent Computing: 7th International Conference, ICIC 2011, Zhengzhou, China, August 11–14. Revised Selected Papers, volume 6838 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24727-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        24727-9.

Huang:2012:AICb

[HGGG12] De-Shuang Huang, Yong Gan, Phalguni Gupta, and M. Michael Gromiha, editors. Advanced Intelligent Computing Theories and Applications. With Aspects of Artificial Intelligence: 7th International Conference, ICIC 2011, Zhengzhou, China, August 11–14, Revised Selected Papers, volume 6839 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25943-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25943-2.

Huang:2012:BIC

[HGPH12] De-Shuang Huang, Yong Gan, Prashan Premaratne, and Kyungsook Han, editors. Bio-Inspired Computing and Applications: 7th International Conference on Intelligent Computing, ICIC 2011, Zhengzhou,China, August 11-14. 2011, Revised Selected Papers, volume 6840 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-24552-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
24552-7.

Harth:2012:CTW

[HK12] Andreas Harth and Nora Koch, editors. Current Trends in Web Engineering: Workshops, Doctoral Symposium, and Tutorials, Held at ICWE 2011, Paphos, Cyprus, June 20–

21. Revised Selected Papers, volume 7059 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27996-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27996-6.

Hameurlain:2012:TLS

[HKW12] Abdelkader Hameurlain, Josef Küng, and Roland Wagner, editors. Transactions on Large-Scale Data- and Knowledge-Centered Systems V, volume 7100 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28147-8. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

                                        pp. URL http://www.
                                        springerlink.com/content/
                                        978-3-642-28147-1.

He:2012:HIS

[HLKX12] Jing He, Xiaohui Liu, Elizabeth A. Krupinski, and Guandong Xu, editors. Health Information Science: First International Conference, HIS 2012, Beijing, China, April 8–10. Proceedings, volume 7231 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29360-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29360-3.

Hao:2012:ECC

[HM12] Jin-Kao Hao and Martin Middendorf, editors. Evolutionary Computation in Combinatorial Optimization: 12th European Conference, EvoCOP 2012, Málaga, Spain, April 11–13. Proceedings, volume 7245 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29123-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29123-4.

Horimoto:2012:ANB

[HNP12]  Katsuhisa  Horimoto,

Masahiko Nakatsui, and Nikolaj Popov, editors. Algebraic and Numeric Biology: 4th International Conference, ANB 2010, Hagenberg, Austria, July 31- August 2, 2010, Revised Selected Papers, volume 6479 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28066-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28066-5.

Ho:2012:AIVb

[Ho12a] Yo-Sung Ho, editor. Advances in Image and Video Technology: 5th Pacific Rim Symposium, PSIVT 2011, Gwangju, South Korea, November 20–23, Proceedings, Part I, volume 7087 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25366-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25366-9.

Ho:2012:AIVa

[Ho12b] Yo-Sung Ho, editor. Advances in Image and Video Technology: 5th Pacific Rim Symposium, PSIVT 2011, Gwangju, South Korea, November 20–23, Proceedings, Part II, volume 7088 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25345-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25345-4.

Herkersdorf:2012:ACS

[HRB12] Andreas Herkersdorf, Kay Römer, and Uwe Brinkschulte, editors. Architecture of Computing Systems — ARCS 2012: 25th International Conference, Munich, Germany, February 28 — March 2, 2012. Proceedings, volume 7179 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28292-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28292-8.

Hadjiantonis:2012:TES

[HS12] Antonis M. Hadjiantonis and Burkhard Stiller, editors. Telecommunication Economics: Selected Results of the COST Action ISO605 Econ@Tel, volume 7216 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30381-1. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

                                        pp. URL http://www.
                                        springerlink.com/content/
                                        978-3-642-30381-4.

Iwamura:2012:CBD

[IS12] Masakazu Iwamura and Faisal Shafait, editors. Camera-Based Document Analysis and Recognition: 4th International Workshop, CB-DAR 2011, Beijing, China, September 22, 2011, Revised Selected Papers, volume 7139 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29363-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29363-4.

Jensen:2012:TPN

[JDK12] Kurt Jensen, Susanna Donatelli, and Jetty Kleijn, editors. Transactions on Petri Nets and Other Models of Concurrency V, volume 6900 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29071-X. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-29071-8.

Joshi:2012:VST

[JMP12] Rajeev Joshi, Peter Müller, and Andreas Podelski, editors. Verified Software: Theories, Tools, Experiments: 4th International Conference, VSTTE 2012, Philadelphia, PA, USA, January 28–29. Proceedings, volume 7152 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27704-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27704-7.

Jonasson:2012:APSa

[Jón12a] Kristján Jónasson, editor. Applied Parallel and Scientific Computing: 10th International Conference, PARA 2010, Reykjavík, Iceland, June 6–9, Revised Selected Papers, Part I, volume 7133 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28150-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28150-1.

Jonasson:2012:APSb

[Jón12b] Kristján Jónasson, editor. Applied Parallel and Scientific Computing: 10th International Conference, PARA 2010, Reykjavík, Iceland, June 6–9, Revised Selected Papers, Part II, volume 7134 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28144-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28144-0.

Juels:2012:RSP

[JP12] Ari Juels and Christof Paar, editors. RFID. Security and Privacy: 7th International Workshop, RFIDSec 2011, Amherst, USA, June 26– 28, Revised Selected Papers, volume 7055 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25285-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25285-3.

Jung:2012:ISA

[JY12] Souhwan Jung and Moti Yung, editors. Information Security Applications: 12th International Workshop, WISA 2011, Jeju Island, Korea, August 22–24. Revised Selected Papers, volume 7115 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27889-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27889-1.

Kannan:2012:DEM

[KA12] Rajkumar Kannan and Frederic Andres, editors. Data Engineering and Management: Second International Conference, ICDEM 2010, Tiruchirappalli, India, July 29– 31. Revised Selected Papers, volume 6411 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27871-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27871-6.

Kotasek:2012:MEM

[KBC^{+}12] Zdenek Kotásek, Jan Bouda, Ivana Cerná, Lukás Sekanina, Tomás Vojnar, and et al, editors. Mathematical and Engineering Methods in Computer Science: 7th International Doctoral Workshop, MEMICS 2011, Lednice, Czech Republic, October 14–16, Revised Selected Papers, volume 7119 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25928-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
25928-9.

Kuipers:2012:SOS

[KH12] Fernando A. Kuipers and Poul E. Heegaard, editors. Self-Organizing Systems: 6th IFIP TC 6 International Workshop, IWSOS 2012, Delft, The Netherlands, March 15–16. Proceedings, volume 7166 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28582-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28582-0.

Kienzle:2012:MSE

[Kie12] Jörg Kienzle, editor.

Models in Software Engineering: Workshops and Symposia at MOD-ELS 2011, Wellington, New Zealand, October 16–21, Reports and Revised Selected Papers, volume 7167 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29644-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29644-4.

Keller:2012:FMC

[KKW12] Rainer Keller, David Kramer, and Jan-Philipp Weiss, editors. Facing the Multicore — Challenge II: Aspects of New Paradigms and Technologies in Parallel Computing, volume 7174 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30396-X. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

                                        pp. URL http://www.
                                        springerlink.com/content/
                                        978-3-642-30396-8.

Kundu:2012:PMI

[KMMP12] Malay K. Kundu, Sushmita Mitra, Debasis Mazumdar, and Sankar K. Pal, editors. Perception and Machine Intelligence: First Indo-Japan Conference, Per-MIn 2012, Kolkata, India, January 12–13. Proceedings, volume 7143 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27386-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27386-5.

Koucheryavy:2012:WWI

[KMMT12] Yevgeni Koucheryavy, Lefteris Mamatas, Ibrahim Matta, and Vassilis Tsaoussidis, editors. Wired/ Wireless Internet Communication: 10th International Conference, WWIC 2012, Santorini, Greece, June 6–8. Proceedings, volume 7277 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30629-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
30629-7.

Kuncak:2012:VMC

[KR12] Viktor Kuncak and Andrey Rybalchenko, editors. Verification, Model Checking, and Abstract Interpretation: 13th International Conference, VMCAI 2012, Philadelphia, PA, USA, January 22–24. Proceedings, volume 7148 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27939-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27939-3.

Khurshid:2012:RVS

[KS12] Sarfraz Khurshid and Koushik Sen, editors. Runtime Verification: Second International Conference, RV 2011, San Francisco, CA, USA, September 27–30, Revised Selected Papers, volume 7186 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29859-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29859-2.

Laud:2012:IST

[Lau12] Peeter Laud, editor. Information Security Technology for Applications: 16th Nordic Conference on Secure IT Systems, NordSec 2011, Tallinn, Estonia, October 26– 28, Revised Selected Papers, volume 7161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29614-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29614-7.

Li:2012:AGP

[LCB12] Ruixuan Li, Jiannong Cao, and Julien Bourgeois, editors. Advances in Grid and Pervasive Computing: 7th International Conference, GPC 2012, Hong Kong, China, May 11–13. Proceedings, volume 7296 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-30766-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
30766-9.

Lukasiewicz:2012:FIK

[LS12] Thomas Lukasiewicz and Attila Sali, editors. Foundations of Information and Knowledge Systems: 7th International Symposium, FoIKS 2012, Kiel, Germany, March 5–9. Proceedings, volume 7153 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28471-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28471-7.

Lones:2012:IPC

[LST^{+}12] Michael A. Lones, Stephen L. Smith, Sarah Teichmann, Felix Naef, James A. Walker, and et al, editors. Information Processing in Cells and Tissues: 9th International Conference, IP-CAT 2012, Cambridge, UK, March 31 — April 2, 2012. Proceedings, volume 7223 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28791-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28791-6.

Lopez:2012:CIP

[LSW12] Javier Lopez, Roberto Setola, and Stephen D. Wolthusen, editors. Critical Infrastructure Protection: Information Infrastructure Models, Analysis, and Defense, volume 7130 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28919-3. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-28919-4.

Lacroix:2012:RDT

[LV12] Zoé Lacroix and María Esther Vidal, editors. Resource Discovery: Third International Workshop, RED 2010, Paris, France, November 5, 2010, Revised Selected Papers, volume 6799 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27391-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27391-9.

MacKinnon:2012:DSS

[Mac12] Lachlan M. MacKinnon, editor. Data Security and Security Data: 27th British National Conference on Databases, BN-COD 27, Dundee, UK, June 29 — July 1, 2010. Revised Selected Papers, volume 6121 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-25703-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25703-2.

Moreno-Diaz:2012:CASa

[MDPQA12a] Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia, editors. Computer Aided Systems Theory — EU-ROCAST 2011: 13th International Conference, Las Palmas de Gran Canaria, Spain, February 6–11, Revised Selected Papers, Part I, volume 6927 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27548-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27548-7.

Moreno-Diaz:2012:CASb

[MDPQA12b] Roberto Moreno-Díaz, Franz Pichler, and Alexis Quesada-Arencibia, editors. Computer Aided Systems Theory — EU-ROCAST 2011: 13th International Conference, Las Palmas de Gran Canaria, Spain, February 6–11, Revised Selected Papers, Part II, volume 6928 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27578-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27578-4.

Meadows:2012:STM

[MFG12] Catherine Meadows and Carmen Fernandez-Gago, editors. Security and Trust Management: 7th International Workshop, STM 2011, Copenhagen, Denmark, June 27–28, Revised Selected Papers, volume 7170 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29962-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29962-9.

Muller:2012:MCB

[MGSM12] Henning Müller, Hayit Greenspan, and Tanveer Syeda-Mahmood, editors. Medical Content-Based Retrieval for Clinical Decision Support: Second MICCAI International Workshop, MCBR-CDS 2011, Toronto, ON, Canada, September 22, 2011, Revised Selected Papers, volume 7075 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28459-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28459-5.

Mossakowski:2012:RTA

[MK12] Till Mossakowski and Hans-Jörg Kreowski, editors. Recent Trends in Algebraic Development Techniques: 20th International Workshop, WADT 2010, Etelsen, Germany, July 1–4, Revised Selected Papers, volume 7137 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28411-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28411-3.

Meyer:2012:ESE

[MN12] Bertrand Meyer and Martin Nordio, editors.

Empirical Software Engineering and Verification: International Summer Schools, LASER 2008-2010, Elba Island, Italy, Revised Tutorial Lectures, volume 7007 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25230-3. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-25230-3.

Modgil:2012:TAF

[MOT12] Sanjay Modgil, Nir Oren, and Francesca Toni, editors. Theorie and Applications of Formal Argumentation: First International Workshop, TAFA 2011. Barcelona, Spain, July 16–17, Revised Selected Papers, volume 7132 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29183-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29183-8.

Modersheim:2012:TSA

[MP12] Sebastian Mödersheim and Catuscia Palamidessi, editors. Theory of Security and Applications: Joint Workshop, TOSCA 2011, Saarbrücken, Germany, March 31 — April 1, 2011, Revised Selected Papers, volume 6993 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27374-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
27374-2.

Martino:2012:WWG

[MPT12] Sergio Di Martino, Adriano Peron, and Taro Tezuka, editors. Web and Wireless Geographical Information Systems: 11th International Symposium, W2GIS 2012, Naples, Italy, April 12– 13. Proceedings, volume 7236 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29246-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29246-0.

Marx:2012:PEC

[MR12] Dániel Marx and Peter Rossmanith, editors. Parameterized and Exact Computation: 6th International Symposium, IPEC 2011, Saarbrücken, Germany, September 6–8. Revised Selected Papers, volume 7112 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28049-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28049-8.

Machado:2012:EBI

[MRC12] Penousal Machado, Juan Romero, and Adrian Carballal, editors. Evolutionary and Biologically Inspired Music, Sound, Art and Design: First International Conference, EvoMUSART 2012, Málaga, Spain, April 11–13. Proceedings, volume 7247 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29141-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29141-8.

Moraglio:2012:GPE

[MSK^{+}12] Alberto Moraglio, Sara Silva, Krzysztof Krawiec, Penousal Machado, and Carlos Cotta, editors. Genetic Programming: 15th European Conference, EuroGP 2012, Málaga, Spain, April 11–13. Proceedings, volume 7244 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29138-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29138-8.

Miri:2012:SAC

[MV12] Ali Miri and Serge Vaudenay, editors. Selected Areas in Cryptography: 18th International Workshop, SAC 2011, Toronto, ON, Canada, August 11–12, Revised Selected Papers, volume 7118 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28495-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28495-3.

Naccache:2012:CST

[Nac12] David Naccache, editor.

Cryptography and Security: From Theory to Applications: Essays Dedicated to Jean-Jacques Quisquater on the Occasion of His 65th Birthday, volume 6805 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28367-5. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-28367-3.

OBoyle:2012:CCI

[O’B12] Michael O’Boyle, editor.

Compiler Construction: 21st International Conference, CC 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7210 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28651-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28651-3.

Ober:2012:SIS

[OO12] Iulian Ober and Ileana Ober, editors. SDL 2011: Integrating System and Software Modeling: 15th International SDL Forum Toulouse, France, July 5–7. Revised Papers, volume 7083 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25263-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25263-1.

Pan:2012:TEV

[PCM^{+}12] Zhigeng Pan, Adrian David Cheok, Wolfgang Müller, Maiga Chang, and Mingmin Zhang, editors. Transactions on Edutainment VII, volume 7145 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29049-3. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

                                        pp. URL http://www.

springerlink.com/content/
978-3-642-29049-7.

Pan:2012:IIDb

[PCN12a] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACI-IDS 2012, Kaohsiung, Taiwan, March 19–21, Proceedings, Part I, volume 7196 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28486-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28486-1.

Pan:2012:IIDc

[PCN12b] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACI-IDS 2012, Kaohsiung, Taiwan, March 19–21, Proceedings, Part II, volume 7197 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28489-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28489-2.

Pan:2012:IIDa

[PCN12c] Jeng-Shyang Pan, Shyi-Ming Chen, and Ngoc Thanh Nguyen, editors. Intelligent Information and Database Systems: 4th Asian Conference, ACI-IDS 2012, Kaohsiung, Taiwan, March 19–21, Proceedings, Part III, volume 7198 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28492-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28492-2.

Picco:2012:WSN

[PH12] Gian Pietro Picco and Wendi Heinzelman, editors. Wireless Sensor Networks: 9th European Conference, EWSN 2012, Trento, Italy, February 15–17. Proceedings, volume 7158 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28168-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28168-6.

Pointcheval:2012:ACE

[PJ12] David Pointcheval and Thomas Johansson, editors. Advances in Cryptology — EUROCRYPT 2012: 31st Annual International Conference on the Theory and Applications of Cryptographic Techniques, Cambridge, UK, April 15–19. Proceedings, volume 7237 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29010-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29010-7.

Petkova-Nikova:2012:PKI

[PNPP12] Svetla Petkova-Nikova, Andreas Pashalidis, and Günther Pernul, editors. Public Key Infrastructures, Services and Applications: 8th European Workshop, EuroPKI 2011, Leuven, Belgium, September 15– 16, Revised Selected Papers, volume 7163 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29803-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29803-5.

Peffers:2012:DSR

[PRK12] Ken Peffers, Marcus Rothenberger, and Bill Kuechler, editors. Design Science Research in Information Systems. Advances in Theory and Practice: 7th International Conference, DESRIST 2012, Las Vegas, NV, USA, May 14–15. Proceedings, volume 7286 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29862-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29862-2.

Pescape:2012:TMA

[PSD12] Antonio Pescapè, Luca Salgarelli, and Xenofontas Dimitropoulos, editors. Traffic Monitoring and Analysis: 4th International Workshop, TMA 2012, Vienna, Austria, March 12, 2012. Proceedings, volume 7189 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28533-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28533-2.

Regnell:2012:REF

[RD12] Björn Regnell and Daniela Damian, editors. Requirements Engineering: Foundation for Software Quality: 18th International Working Conference, REFSQ 2012, Essen, Germany, March 19–22. Proceedings, volume 7195 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28713-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28713-8.

Rahman:2012:WAC

[RiN12] Md. Saidur Rahman and Shin ichi Nakano, editors. WALCOM: Algorithms and Computation: 6th International Workshop, WAL-COM 2012, Dhaka, Bangladesh, February 15–17. Proceedings, volume 7157 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28075-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28075-7.

Rautiainen:2012:GPC

[RKM^{+}12] Mika Rautiainen, Timo Korhonen, Edward Mutafungwa, Eila Ovaska, Artem Katasonov, and et al, editors. Grid and Pervasive Computing Workshops: International Workshops, S3E, HWTS, Doctoral Colloquium, Held in Conjunction with GPC 2011, Oulu, Finland, May 11– 13. Revised Selected Papers, volume 7096 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27915-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27915-7.

Rutkowski:2012:AISa

[RKS^{+}12a] Leszek Rutkowski, Marcin Korytkowski, Rafal Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and et al, editors. Artificial Intelligence and Soft Computing: 11th International Conference, ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012, Proceedings, Part I, volume 7267 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29346-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29346-7.

Rutkowski:2012:AISb

[RKS^{+}12b] Leszek Rutkowski, Marcin Korytkowski, Rafal Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and et al, editors. Artificial Intelligence and Soft Computing: 11th International Conference, ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012, Proceedings, Part II, volume 7268 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29349-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.

                                        com/content/978-3-642-
                                        29349-8.

Rutkowski:2012:SEC

[RKS^{+}12c] Leszek Rutkowski, Marcin Korytkowski, Rafal Scherer, Ryszard Tadeusiewicz, Lotfi A. Zadeh, and et al, editors. Swarm and Evolutionary Computation: International Symposia, SIDE 2012 and EC 2012, Held in Conjunction with ICAISC 2012, Zakopane, Poland, April 29–May 3, 2012. Proceedings, volume 7269 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29352-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29352-8.

Ramanujam:2012:DCI

[RR12] R. Ramanujam and Srini Ramaswamy, editors. Distributed Computing and Internet Technology: 8th International Conference, ICDCIT 2012, Bhubaneswar, India, February 2–4. Proceedings, volume 7154 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28072-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28072-6.

Ryan:2012:ISP

[RSW12] Mark D. Ryan, Ben Smyth, and Guilin Wang, editors. Information Security Practice and Experience: 8th International Conference, IS-PEC 2012, Hangzhou, China, April 9–12. Proceedings, volume 7232 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29100-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29100-5.

Riano:2012:KRH

[RtTM12] David Riaño, Annette ten Teije, and Silvia Miksch, editors. Knowledge Representation for Health-Care: AIME 2011 Workshop KR4HC 2011, Bled, Slovenia, July 2– 6. Revised Selected Papers, volume 6924 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27696-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27696-5.

Russo:2012:PAD

[RZ12] Claudio Russo and Neng-Fa Zhou, editors. Practical Aspects of Declarative Languages: 14th International Symposium, PADL 2012, Philadelphia, PA, USA, January 23–24. Proceedings, volume 7149 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-27693-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27693-4.

Sloane:2012:SLE

[SA12] Anthony Sloane and Uwe Aßmann, editors. Software Language Engineering: 4th International Conference, SLE 2011, Braga, Portugal, July 3– 4, Revised Selected Papers, volume 6940 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28829-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28829-6.

Schmitt:2012:MME

[Sch12] Jens B. Schmitt, editor. Measurement, Modelling, and Evaluation of Computing Systems and Dependability and Fault Tolerance: 16th International GI/ITG Conference, MMB & DFT 2012, Kaiserslautern, Germany, March 19–21. Proceedings, volume 7201 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28539-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28539-4.

Seidl:2012:PLS

[Sei12] Helmut Seidl, editor.

Programming Languages and Systems: 21st European Symposium on Programming, ESOP 2012, Held as Part of the European Joint Conferences on Theory and Practice of Software, ETAPS 2012, Tallinn, Estonia, March 24 — April 1, 2012. Proceedings, volume 7211 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28868-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28868-5.

Sanner:2012:RAR

[SH12a] Scott Sanner and Marcus Hutter, editors. Recent Advances in Reinforcement Learning: 9th European Workshop, EWRL 2011, Athens, Greece, September 9– 11, Revised Selected Papers, volume 7188 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29945-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29945-2.

Schindler:2012:CSC

[SH12b] Werner Schindler and Sorin A. Huss, editors. Constructive Side-Channel Analysis and Secure Design: Third International Workshop, COSADE 2012, Darmstadt, Germany, May 3–4. Proceedings, volume 7275 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29911-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29911-7.

Shi:2012:TDH

[Shi12] Yun Q. Shi, editor.

Transactions on Data Hiding and Multimedia Security VII, volume 7110 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28692-5. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-28692-6.

Snoeyink:2012:FAA

[SLSW12] Jack Snoeyink, Pinyan Lu, Kaile Su, and Lusheng Wang, editors. Frontiers in Algorithmics and Algorithmic Aspects in Information and Management: Joint International Conference, FAW-AAIM 2012, Beijing, China, May 14–16. Proceedings, volume 7285 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29699-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29699-4.

Schoeffmann:2012:AMM

[SMH^{+}12] Klaus Schoeffmann, Bernard Merialdo, Alexander G. Hauptmann, Chong-Wah Ngo, Yiannis Andreopoulos, and et al, editors. Advances in Multimedia Modeling: 18th International Conference, MMM 2012, Klagenfurt, Austria, January 4–6. Proceedings, volume 7131 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-27354-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        27354-4.

Solis-Oba:2012:AOA

[SOP12] Roberto Solis-Oba and Giuseppe Persiano, editors. Approximation and Online Algorithms: 9th International Workshop, WAOA 2011, Saarbrücken, Germany, September 8–9, Revised Selected Papers, volume 7164 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29115-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29115-9.

Sakama:2012:DAL

[SSVW12] Chiaki Sakama, Sebastian Sardina, Wamberto Vasconcelos, and Michael Winikoff, editors. Declarative Agent Languages and Technologies IX: 9th International Workshop, DALT 2011, Taipei, Taiwan, May 3, 2011, Revised Selected and Invited Papers, volume 7169 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29112-0. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29112-8.

Szmuc:2012:ASE

[SSZ12] Tomasz Szmuc, Marcin Szpyrka, and Jaroslav Zendulka, editors. Advances in Software Engineering Techniques: 4th IFIP TC 2 Central and East European Conference on Software Engineering Techniques, CEE-SET 2009, Krakow, Poland, October 12–14. Revised Selected Papers, volume 7054 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28037-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28037-5.

Schrijvers:2012:FLP

[ST12a] Tom Schrijvers and Peter Thiemann, editors. Functional and Logic Programming: 11th International Symposium, FLOPS 2012, Kobe, Japan, May 23–25. Proceedings, volume 7294 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29821-4. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29821-9.

Schwenker:2012:PSL

[ST12b]  Friedhelm  Schwenker

and Edmondo Trentin, editors. Partially Supervised Learning: First IAPR TC3 Workshop, PSL 2011, Ulm, Germany, September 15– 16, Revised Selected Papers, volume 7081 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28257-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28257-7.

Sheng:2012:WTA

[SWJX12] Quan Z. Sheng, Guoren Wang, Christian S. Jensen, and Guandong Xu, editors. Web Technologies and Applications: 14th Asia-Pacific Web Conference, AP-Web 2012, Kunming, China, April 11–13. Proceedings, volume 7235 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29252-6. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29252-1.

Tan:2012:AKDb

[TCHB12a] Pang-Ning Tan, Sanjay Chawla, Chin Kuan Ho, and James Bailey, editors. Advances in Knowledge Discovery and Data Mining: 16th Pacific-Asia Conference, PAKDD 2012, Kuala Lumpur, Malaysia, May 29 — June 1, 2012, Proceedings, Part II, volume 7302 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30219-X. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
30219-0.

Tan:2012:AKDa

[TCHB12b] Pang-Ning Tan, Sanjay Chawla, Chin Kuan Ho, and James Bailey, editors. Advances in Knowledge Discovery and Data Mining: 16th Pacific-Asia Conference, PAKDD 2012, Kuala Lumpur, Malaysia, May 29–June 1, 2012, Proceedings, Part I, volume 7301 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-30216-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
30216-9.

Theis:2012:LVA

[TCYZ12] Fabian Theis, Andrzej Cichocki, Arie Yeredor, and Michael Zibulevsky, editors. Latent Variable Analysis and Signal Separation: 10th International Conference, LVA/ ICA 2012, Tel Aviv, Israel, March 12–15. Proceedings, volume 7191 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28550-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28550-9.

Thilagam:2012:ACN

[TPCB12] P. Santhi Thilagam, Alwyn Roshan Pais, K. Chandrasekaran, and N. Balakrishnan, editors. Advanced Computing, Networking and Security: International Conference, ADCONS 2011, Surathkal, India, December 16–18, Revised Selected Papers, volume 7135 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29279-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29279-8.

Taft:2012:PAM

[TR12] Nina Taft and Fabio Ricciato, editors. Passive and Active Measurement: 13th International Conference, PAM 2012, Vienna, Austria, March 12-14th, 2012. Proceedings, volume 7192 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28536-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28536-3.

Vanmechelen:2012:EGC

[VAR12] Kurt Vanmechelen, Jörn Altmann, and Omer F. Rana, editors. Economics of Grids, Clouds, Systems, and Services: 8th International Workshop, GECON 2011, Paphos, Cyprus, December 5, 2011, Revised Selected Papers, volume 7150 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28674-7. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28674-2.

vanEijck:2012:GAS

[vEV12] Jan van Eijck and Rineke Verbrugge, editors. Games, Actions and Social Software: Multidisciplinary Aspects, volume 7010 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29327-1. ISSN 0302-9743 (print), 1611-3349 (electronic). ????

pp. URL http://www.
springerlink.com/content/
978-3-642-29327-6.

Vrancx:2012:ALA

[VKG12] Peter Vrancx, Matthew Knudson, and Marek Grzes, editors. Adaptive and Learning Agents: International Workshop, ALA 2011, Held at AA-MAS 2011, Taipei, Taiwan, May 2, 2011, Revised Selected Papers, volume 7113 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28498-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28498-4.

vanKreveld:2012:GDI

[vKS12] Marc van Kreveld and Bettina Speckmann, editors. Graph Drawing: 19th International Symposium, GD 2011, Eindhoven, The Netherlands, September 21– 23, Revised Selected Papers, volume 7034 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-25877-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        25877-0.

Vinel:2012:CTV

[VMB^{+}12] Alexey Vinel, Rashid Mehmood, Marion Berbineau, Cristina Rico Garcia, Chung-Ming Huang, and et al, editors. Communication Technologies for Vehicles: 4th International Workshop, Nets4Cars/Nets4Trains 2012, Vilnius, Lithuania, April 25–27. Proceedings, volume 7266 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29666-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29666-6.

Varbanescu:2012:CAI

[VMvN12] Ana Lucia Varbanescu, Anca Molnos, and Rob van Nieuwpoort, editors. Computer Architecture: ISCA 2010 International Workshops A4MMC, AMAS-BT, EAMA, WEED, WIOSCA, Saint-Malo, France, June 19–23, Revised Selected Papers, volume 6161 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-24321-5. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
24321-9.

Villatoro:2012:MAB

[VSMS12] Daniel Villatoro, Jordi Sabater-Mir, and Jaime Simão Sichman, editors. Multi-Agent-Based Simulation XII: International Workshop, MABS 2011, Taipei, Taiwan, May 2–6, Revised Selected Papers, volume 7124 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-28399-3. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28399-4.

Vos:2012:RCT

[VW12] Alexis De Vos and Robert Wille, editors. Reversible Computation: Third International Workshop, RC 2011, Gent, Belgium, July 4–5. Revised Papers, volume 7165 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CODEN LNCSD9. ISBN 3-642-29516-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://www.springerlink.
                                        com/content/978-3-642-
                                        29516-4.

Wang:2012:WAI

[WJL^{+}12] Liwei Wang, Jingjue Jiang, Jiaheng Lu, Liang Hong, and Bin Liu, editors. Web-Age Information Management: WAIM 2011 International Workshops: WGIM 2011, XMLDM 2011, SNA 2011, Wuhan, China, September 14– 16, Revised Selected Papers, volume 7142 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28634-8. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
28634-6.

Wang:2012:WTA

[WZH^{+}12] Hua Wang, Lei Zou, Guangyan Huang, Jing He, Chaoyi Pang, and et al, editors. Web Technologies and Applications: APWeb 2012 International Workshops: SenDe, IDP, IEKB, MBC, Kunming, China, April 11–13. Proceedings, volume 7234 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29425-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29425-9.

Yang:2012:SCB

[YGE12] Shanchieh Jay Yang, Ariel M. Greenberg, and Mica Endsley, editors. Social Computing, Behavioral — Cultural Modeling and Prediction: 5th International Conference, SBP 2012, College Park, MD, USA, April 3–5. Proceedings, volume 7227 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29046-9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        29046-6.

Yoshida:2012:AIC

[YSL12] Hiroyuki Yoshida, Georgios Sakas, and Marius George Linguraru, editors. Abdominal Imaging. Computational and Clinical Applications: Third International Workshop, Held in Conjunction with MIC-CAI 2011, Toronto, ON, Canada, September 18, 2011, Revised Selected Papers, volume 7029 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-28556-2. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //www.springerlink.
                                        com/content/978-3-642-
                                        28556-1.

Yu:2012:DSA

[YYH^{+}12] Hwanjo Yu, Ge Yu, Wynne Hsu, Yang-Sae Moon, Rainer Unland, and et al, editors. Database Systems for Advanced Applications: 17th International Conference, DAS-FAA 2012, International Workshops: FlashDB, ITEMS, SNSM, SIM^{3}, DQDI, Busan, South Korea, April 15–19. Proceedings, volume 7240 of Lecture Notes in Computer Science. Springer-Verlag Inc., New York, NY, USA, 2012. CO-DEN LNCSD9. ISBN 3-642-29022-1. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//www.springerlink.
com/content/978-3-642-
29022-0.